OpenStudy (anonymous):

A hat contains 24 names, 13 of which are female. If three names are randomly drawn from the hat, what is the probability that at least one male name is drawn?

OpenStudy (anonymous):

13 female names, 11 male names. For the above probability NOT to happen, no male names must be drawn, ALL female, which is 13/24*12/23*11/22 = 13/92 Therefore, the probability it WILL happen is 1-13/92 = 79/92

OpenStudy (anonymous):

my options are 0.918, 0.859, 0.082, or 0.141